OpenStudy (anonymous):

Ok, can someone help me understand this question? Coach Johnson is 10 years older than Tim. When Tim started football 3 years ago, Coach Johnson was twice his age. How old is Tim now? Or at least help me with a formula to figure it out, I am just stumped.

OpenStudy (anonymous):

Let's call Coach Johnson's age x, and Tim's age y. So you know that \[x+10=y\] You also know that 3 years ago, Coach Johnson's age would be x-3 and Tim's would be y-3. If Coach Johnson was twice his age back then, then\[2(x-3)=y-3\] You have a system of two equations with two unknowns, now solv!

OpenStudy (anonymous):

Solve for x and y that is.

OpenStudy (anonymous):

Oh wow, lol I feel silly, thank you for the formula.

OpenStudy (anonymous):

Woops, I made a mistake! The 2nd equation should be \[(x-3)=2(y-3)\] Hope that helps.